# WarmSwap: Sharing Dependencies for Accelerating Cold Starts in Serverless Functions

### v1

@@ Abstract
This work presents WarmSwap, a novel provider-side cold-start optimization for serverless computing. This optimization reduces cold-start time when booting and loading dependencies at runtime inside a function container. Previous approaches to the optimization of cold starts tend to fall into two categories: optimizing the infrastructure of serverless computing to benefit all serverless functions; or function-specific tuning for individual serverless functions. In contrast, WarmSwap offers a broad middle ground, which optimizes entire categories of serverless functions. WarmSwap eliminates the need to initialize middleware or software dependencies when launching a new serverless container, by migrating a pre-initialized live dependency image to the new function instance. WarmSwap respects the provider’s cache constraints, as a single pre-warmed dependency image in the cache is shared among all serverless functions requiring that software dependency image. WarmSwap has been tested on seven representative functions from FunctionBench. The functions are chosen to compare with previous work. In those tests, WarmSwap accelerates cold-start executions for those serverless functions with large dependency requirements by a factor ranging from 1.2 to 2.2.
1 Introduction
@@
Serverless computing is a flexible and popular cloud computing methodology that provides users with a simpler and cheaper way to deploy applications. It has quickly gained attention across various cloud providers, including AWS [3], Azure [19], Google[11], and Alibaba [2]. For the user, serverless computing offers robust auto-scaling and an economical pay-as-you-go model, in contrast to reserving an instance for a fixed time. Users only need to submit their code as a serverless function in the cloud and later trigger the function’s execution based on a request or event. Finally, the user’s code is executed in instances (e.g., a virtual machine or a Docker container) provided by cloud providers.
@@
However, many of the financial benefits are undermined by the cold-start problem. In serverless computing, the cloud provider loads user function into the serverless instance only upon user invocation, forming the basis of the “pay-as-you-go” pricing model. Typically, the cloud retains the serverless instance in memory for a certain period (i.e., during a keep-alive time).
If an invocation occurs when no instance is alive in the system, the function experiences a cold start, which is significantly slower due to instance initialization. If the next invocation happens within the keep-alive time, the function experiences a warm start, during which the provider does not need to load and initialize a new instance. Cold starts can also occur during spikes in invocations when the system requires additional instances of a given serverless function. However, as reported by Azure [22], this affects only a small percentage of applications (less than 1%).
Given the significance of the cold-start problem [10], recent years have witnessed a surge in related research efforts from both academia and industry [6, 8, 21, 23, 26, 20, 15, 24]. To address the overhead of cold start, these researchers propose various system designs that leverage techniques such as caching, remote memory, and checkpoint/restore. Their ultimate goal is to achieve a faster cold start by accelerating or bypassing certain initialization phases.
Optimizations for serverless cold start can be classified into function-specific tuning and infrastructure optimization based on their targets. Function-specific tuning focuses on individual functions, such as creating and caching pre-initialized images [23, 8, 6], generating function seeds [26], and recording and pre-fetching hot pages [25]. Infrastructure optimization typically targets the entire system or shareable resources to benefit different functions. Examples include building a more efficient storage or caching system [5, 1], maintaining pre-initialized Zygote containers [21, 16], and sharing idle containers across different functions [14].
Function-specific tuning typically results in very fast cold starts because it can manage every detail of the initialization phase for a specific function. However, implementing these optimizations for all serverless functions on the cloud can incur substantial costs for the cloud provider. In contrast, infrastructure optimizations have lower costs for large-scale cloud providers because they naturally benefit all serverless functions running on the cloud. Furthermore, infrastructure optimizations are easier to implement in real-world cloud environments because they only require involvement from the cloud providers.
@@
Figure 1: The relationship of invocation rate $\lambda$ versus number of expected cold starts with keep-alive time $T=15$ minutes and the function distribution coming from Azure [22]. Each serverless function was assigned to a bucket $[x,x+0.001]$, where the $\lambda$ of the serverless function was between $x$ and $x+0.001$ invocations per minute. The y-value for a bucket is the number of functions that fall into the given bucket. The y-values for each bucket were normalized so that the sum of the values for the buckets adds to one.
In practice, production cloud providers tend to emphasize more infrastructure optimization, due to the skewed workload in the real world, as illustrated in Figure 1. Function-specific tuning is suitable for functions with invocation rates around $1/T$, where a significant number of expected cold starts occur to each function. However, according to an Azure Function dataset [22], 50% of functions experience fewer than $1.4$ cold starts per day, assuming a keep-alive time of 15 minutes. Therefore, it is costly and unnecessary to maintain pre-initialized images or record hot pages in RAM for all serverless functions.
Although some infrastructure optimization for serverless container caching and launching have been successfully implemented in production [1, 5], there remains a lack of infrastructure optimization for loading software dependencies. For instance, AWS Lambda can launch cold starts for function containers up to 10 GB in size within 50 ms [5], but it still initializes software dependencies from scratch (see experiments in Section 2.3). The time required to load software dependencies within the container can sometimes exceed the execution time [21]. Recently, this initialization time has been exacerbated by the development of AI and large language models, as AI-related functions often require larger and more complex dependencies.
The existing work targets reducing dependency-loading latency, but deploying those solutions in real-world serverless systems still presents challenges. SOCK [21] and FlashCube [16] have proposed new container frameworks that aim to bypass container launch and dependency-loading phases by preserving Zygote containers and container components on serverless worker nodes. However, preserving pre-initialized containers or their components on every serverless worker node, or remotely migrating the entire container, is costly. It also restricts user flexibility to use their own custom containers, a feature commonly supported by large serverless providers like AWS [3].
Finally, this work is complementary to previous work, and can effectively coexist with previous approaches. The function-specific tuning like catalyzer [8] or vHive [25] optimize the frequently used serverless functions (i.e. around $1/T$ in Figure 1). This work optimizes the “long tail” of the distribution of buckets for serverless functions (see caption of Figure 1), as documented by Azure [22]. And finally, infrastructure optimizations benefit [5, 1] all serverless functions.
The WarmSwap Approach:
@@
In this paper, we present WarmSwap, targeting the dependency loading overhead during cold starts. In WarmSwap, the provider maintains a pool of live dependency images in RAM. A live dependency image is an image stored in memory that represents a checkpoint of the software dependencies. Typically, these images contain popular middleware combinations, such as Python with NumPy, Python with Torch, or Python with NumPy and Torch.
WarmSwap is designed to benefit serverless providers by also optimizing the massive number of less frequently invoked functions (the “long tail” of the distribution), which is usually not practical in a function-specific approach. WarmSwap brings several benefits. First, providers only need to maintain a limited number of live dependency images in memory, which can be shared by a wide variety of serverless functions that utilize that middleware. The pre-warming cost for the provider depends solely on the number of live dependency images, not on the number of serverless functions employed by multiple users. However, the benefit of pre-initialization is shared among all serverless functions using the middleware. Second, WarmSwap respects isolation for serverless functions. This is because the provider primarily transfers pages from public dependency images, which do not include user source code.
By using WarmSwap, the provider can accelerate the cold start by “pre-warming” a popular dependency pool in either the local worker nodes or a remote central node. When creating new serverless instances for cold starts, WarmSwap can first check whether a necessary software dependency already exists in the dependency image pool. If so, WarmSwap conducts a live migration of the required dependency image from the pool to new serverless containers. This saves time otherwise spent on booting runtimes and loading software dependencies.
Furthermore, the WarmSwap approach is practical because it focuses on widely used, public software dependencies. A study [21] reveals that 36% of “import” in 876,000 Python projects are from just 20 packages. Therefore, providers can concentrate on maintaining a relatively small number of dependency images across numerous serverless functions used by all users.
@@
An experimental implementation of WarmSwap was evaluated by running offline AWS Lambda containers on AWS EC2 nodes. In our tests, loading libraries such as “sklearn+pandas”, “numpy+keras”, and “numpy+torch” through WarmSwap was up to 2.2x, 3.2x, and 2.5x faster, respectively, compared to traditional AWS Lambda startup processes. Furthermore, WarmSwap can save 85.4% of memory space compared to the function-specific approach [23] when warming up 10 different serverless functions sharing the same dependency. Therefore, for functions sharing a common dependency, WarmSwap represents a new type of infrastructure optimization using pre-warmed dependency images, offering several advantages over previous approaches in the literature.
@@
    This paper proposes WarmSwap, a new infrastructure optimization targeting the loading of a software dependency to accelerate serverless cold starts, benefiting the many less frequently invoked serverless functions as well.
@@
    WarmSwap is a cost-effective solution for dependency-loading, as the provider stores middleware only once in a separate dependency image, independent of the entire container. A single pre-warmed dependency image can be shared between different serverless functions and worker nodes.
@@
    This work is targeted to benefit the real-world distributions of serverless function invocations with their “long tail”, as documented by an Azure study (see Figure 1).
The rest of this paper is organized as: Section 2 presents the background and motivation of our paper; Section 3 presents the design and implementation details of WarmSwap; In Section  4, we evaluate the performance of WarmSwap; We then discuss related work in Section 5, and conclude in Section 6.
2 Background and Motivation
This section introduces the background for serverless computing and the issue of cold starts in Section 2.1. Section 2.2 motivates the need for WarmSwap by analyzing the relationship between cold starts and function invocation patterns in Azure. Section 2.3 then discusses an existing performance bottleneck in a commercial cloud platform such as AWS Lambda.
2.1 Cold Start in Serverless Computing
@@
Figure 2: The process of serverless cold starts in the current production cloud: Different optimization methods may involve one or more steps in the process.
In Figure 2, while the specifics may differ across platforms, most serverless architectures typically consist of two core components: a Worker Manager and a Resource Orchestration unit. In response to a function invocation from the front end, a request for capacity is sent to the Worker Manager, which acts as a stateful load balancer. The Worker Manager continuously monitors the capacity of each unique function on the platform. If sufficient capacity exists (e.g., idle function instances are available), the Worker Manager directs the front end to route the request to an idle instance. Conversely, if resources are unavailable (triggering a cold start), it instructs Resource Orchestration to provision a new function instance with adequate CPU and RAM.
Recall also that the user has uploaded the container function to the provider’s Container Registry. Next, the provider will download the user container from the Container Registry and start it. The container initiates the serverless runtime to boot the entire serverless function. Then, the serverless runtime starts a specific function runtime (e.g., Python), and the program loads the necessary software dependencies (e.g., NumPy) and the user-specific handler code. At this point, the new function instance is ready to handle user requests.
To minimize the cold-start overhead, one needs to skip or expedite one or more steps in the cold-start optimization. When designing a function-specific tuning for the cold start, one can consider all the steps specific to a serverless function. Previous works have designed end-to-end, function-specific tuning for targeted serverless functions and achieved substantial reductions in cold-start latency. However, when performing infrastructure optimization for the entire system, one can only optimize steps that are identical or similar across all serverless functions, such as container storage and loading.
In this paper, we argue that runtime booting and dependency-loading qualify for infrastructure optimization because they are common processes that can benefit multiple serverless functions. Thus, we propose WarmSwap as a novel infrastructure optimization aimed at reducing the time required for runtime booting and dependency-loading.
2.2 Impact of Invocation Patterns on Cold Starts
In this section, we aim to address two questions:
  1. 1)
    How frequently does a cold start occur for a serverless function?
  2. 2)
    How does a skewed workload impact the optimization choices of the cloud provider?
@@ 2.2 Impact of Invocation Patterns on Cold Starts
To address question 1, we model the occurrence of cold starts using an exponential distribution. The frequency of cold starts for a serverless function depends on its invocation pattern. For extreme cases, a function that is never invoked or invoked very frequently (e.g., every second) would not experience cold starts. Let $\lambda$ denote the invocation rate (calls per minute), and $T$ denote the keep-alive time. The probability of no invocation occurring during the keep-alive time follows an exponential distribution:
@@
Here, $P_{no\_inv}$ represents the probability that no invocation happens during the keep-alive timeout, leading to a potential cold start.
@@
Figure 1 shows the relation of $E_{cs}$ versus $\lambda$ when $T=10\;min$ and $D=1440\;min$ (a day). Only functions with invocation rates around $1/T$ have a large number of expected cold starts, functions that are less or more frequently invoked have fewer cold starts. Cloud providers like Azure [19] and AWS [3] configure dynamic keep-alive time from $5$ to $30$ minutes. The different keep-alive time changes the position of the maximum point, but the shapes are similar.
To answer question two, we draw the density histogram of the actual function distribution in the same Figure 1. For function-specific tuning, assuming the cost of optimizing a function is $c$ and the benefits of an optimized cold start are $w$. Functions that are worth function-specific tuning must satisfy the following:
@@
However, from the histogram, the serverless workload in production can be extremely skewed. Over $50\%$ of functions have very a low invocation rate (less than 0.001 calls per minute) and $E_{cs}$, which experience less than $1.4$ times of cold starts. Thus, those functions are hard to qualify equation 3, which means function-specific tuning is not economic for them.
For those functions whose invocation rate cannot qualify equation 3, only infrastructure optimization can help to reduce their cold-start overhead. One successful optimization example from the provider is AWS Lambda, which published its optimization of function creation and container loading in a recent paper [5]. The design of AWS lambda assumes that serverless instances are highly cacheable, in which most scaling is driven by a smaller number of images, and the images are based on common base layers. With a three-tier caching system and container deduplication, AWS can cold start a 10 GB serverless container within 50 ms. Thus, the loading and booting of the container are lightweight in current commercial serverless systems.
In this paper, we hope WarmSwap can further extend the infrastructure optimization for the benefit of the many serverless functions that can not apply function-specific tuning.
2.3 Performance Analysis of the Cold Start for Current AWS Lambda Functions
This section analyzes the performance of cold-start overhead in a production cloud on AWS [1, 5]. The goal is to analyze which components of the current AWS Lambda architecture still have room for optimization. None of the experiments in this section refer to WarmSwap.
For analysis, we built serverless function containers to create AWS Lambda functions, and then invoke them by a script running on an AWS EC2 node. All Lambda functions are configured with a 10 GB memory limitation to maximize the CPU resources. To collect cold-start times, the script needs to wait until the end of keep-alive time between invocations of the same function. There are several timers in the analysis script and the function containers to track the time of each cold-start process. The entire time for the cold start is calculated from the script that sends the request to the serverless function and receives the result.
Note that the caching system of AWS Lambda would bring a bias to a naive experimental design. For example, if the user uploads a new function container, the first cold start will be extremely slow as compared to the following ones. Moreover, if the functions are frequently invoked, AWS upgrades the function to a different cache that will improve its cold starts. (See “AWS three-tier caching” in [5].) The improved caching scheme continues to be used even when the function is deleted and then recreated. In order to achieve reproducible numerical results, we upload a container and then invoke the serverless function many times, until the cold-start results are stable.
Serverless Functions Container Size (MB) Major Dependencies
helloworld 180.00 None
json_dumps_load 180.00 urllib, json
pyase 180.11 pyaes, string
chameleon 180.42 chameleon
lr_serving 379.52 sklearn, pandas
cnn_serving 1386.73 numpy, keras
rnn_serving 5602.46 numpy, torch
Table 1: Serverless functions adopted from FunctionBench [12]
Table 1 presents serverless functions from a representative serverless benchmark suite called FunctionBench [12, 13]. We have chosen FunctionBench, for ease of comparison with previous work [25]. All these functions are written in Python and they use some representative Python packages like “numpy”, “torch”, “keras”, and “sklearn”. Table 1 lists both packages and the size of specific function container images. Installing large packages also significantly increases the image size of function containers. Section 4.1 describes more details of these functions.

### v3

# Boston, MA, USA

@@ Abstract
This work presents HotSwap, a novel provider-side cold-start optimization for serverless computing. This optimization reduces cold-start time when booting and loading dependencies at runtime inside a function container. Previous research has extensively focused on reducing cold-start latency for specific functions. However, little attention has been given to skewed production workloads. In such cases, cross-function optimization becomes essential. Without cross-function optimization, a cloud provider is left with two equally poor options: (i) Either the cloud provider gives up optimization for each function in the long tail (which is slow); or (ii) the cloud provider applies function-specific optimizations (e.g., cache function images) to every function in the long tail (which violates the vendor’s cache constraints). HotSwap demonstrates cross-function optimization using a novel pre-warming strategy. In this strategy, a pre-initialized live dependency image is migrated to the new function instance. At the same time, HotSwap respects the provider’s cache constraints, because a single pre-warmed dependency image in the cache can be shared among all serverless functions that require that image. HotSwap has been tested on seven representative functions from FunctionBench. In those tests, HotSwap accelerates dependency loading for those serverless functions with large dependency requirements by a factor ranging from 2.2 to 3.2. Simulation experiments using Azure traces indicate that HotSwap can save 88% of space, compared with a previous function-specific method, PreBaking, when sharing a dependency image among ten different functions.
Index Terms:
Cloud Computing, FaaS, Serverless Computing, Cold Start, Dependency Pre-Warming
I Introduction
@@
Serverless computing is a flexible and popular cloud computing methodology that provides users with a simpler and cheaper way to deploy applications. A common example is Function-as-a-Service (FaaS), which enables event-driven execution of individual functions. It has quickly gained attention across various cloud providers, including AWS [1], Azure [2], Google[3], and Alibaba [4].
@@
However, many of the financial benefits are undermined by the cold-start problem. In FaaS, the cloud provider loads user function into the serverless instance only upon user invocation. After an invocation, the cloud retains the serverless instance in memory for a certain period (i.e., during a keep-alive time). If an invocation occurs when no instance is alive in the system, the function experiences a cold start, which is significantly slower due to instance initialization. Otherwise, the function experiences a faster warm start with no initialization phase. Due to the significance and challenges of the cold start problem, it has been extensively studied by systems researchers [5, 6, 7].
A naive solution to avoid the cold start is to always keep the function instance alive, ensuring that cold starts never occur. Researchers have further developed approaches to intelligently adjust the keep-alive duration based on invocation patterns [8, 9] or by predicting invocation arrivals and pre-loading functions in advance [10, 8]. However, most user request arrivals are hard to predict [11, 12].
Keeping idle functions alive wastes resources, prompting research into accelerating cold starts instead. A common approach is to pre-warm functions using checkpoint images or caches [13, 14, 15, 16], which bypass parts of the initialization process. While effective, these methods increase memory usage, straining limited cache resources. Recent work refines this by pre-fetching hot pages [17] and generating function seeds [18], but further optimization is needed to reduce overhead and improve performance.
Experiments from a recent state-of-the-art work show that these methods improve the cold-start tail latency by 89% and require only 29MB of memory per machine to pre-warm a serverless function [18]. These methods optimize cold starts by analyzing function-specific initialization bottlenecks. However, they are not lightweight, often requiring deep analysis or specialized OS and hardware support [18], such as RDMA.
Nevertheless, according to Azure Function traces [11], 50% of functions experience fewer than $1.4$ cold starts per day, assuming a keep-alive time of 15 minutes. Therefore, maintaining pre-initialized images or recording hot pages in RAM for every serverless function in a workload like Azure [11] is costly and unnecessary.
Implementing function-specific optimizations for all serverless functions is impractical in production due to skewed workloads, where many functions experience few cold starts (see Figure 1). For functions with invocation rates far from $1/T$, the cost of applying prior methods is not economical, where $T$ is the keep-alive time.
A more practical approach, albeit with other limitations, is cross-function pre-warming, where shared resources are preserved across functions [19, 20, 21]. For instance, SOCK [19] and FlashCube [21] reduce startup latency respectively by retaining Zygote containers or by retaining shared components on worker nodes. However, maintaining or migrating such pre-initialized containers remains costly at scale.
@@
Figure 1: The relationship between invocation rate $\lambda$ and expected cold starts with keep-alive time $T=15$ minutes, using function distribution from Azure [11]. Each serverless function was assigned to a bucket $[x,x+0.001]$, where the $\lambda$ of the serverless function was between $x$ and $x+0.001$ invocations per minute. The y-value for a bucket is the normalized number of functions in that bucket.
Despite extensive research on the cold start problem, key challenges remain and hinder large-scale deployment in production. In particular, there is a lack of lightweight, cost-efficient, cross-function optimizations that target the majority of infrequently invoked functions (see the caption of Figure 1). This paper aims to enable cross-function optimizations in the dependency-loading phase, a major bottleneck in current production cloud environments (see Section II-C), while keeping overhead minimal. Thus, we need only apply cross-function optimization to the frequently used dependencies instead of entire images, thereby improving scalability.
The HotSwap Approach:
@@
In this paper, we present HotSwap, a cross-function solution targeting the dependency-loading overhead during cold starts. HotSwap enables the provider to maintain a pool of live dependency images in RAM. A live dependency image is a memory-resident checkpoint representing software dependencies. Typically, these images include popular software dependency combinations, such as NumPy and Torch.
HotSwap is designed to benefit serverless providers by also optimizing the massive number of less frequently invoked functions (the “long tail” of the distribution), which is usually not practical in a function-specific approach. HotSwap brings an important benefit. Providers only need to maintain a limited number of live dependency images in memory, which can be shared across a wide range of serverless functions using that dependency. Therefore, for functions sharing a common dependency, HotSwap represents a new type of cross-function optimization using pre-warmed dependency images, offering several advantages over previous approaches in the literature. Supporting this, a study [19] analyzed 876,000 Python projects and found that just 20 packages accounted for 36% of all ‘import’ statements, indicating that most functions rely on a small set of common dependencies.
@@
A prototype of HotSwap was evaluated by running offline AWS Lambda containers on AWS EC2 nodes. The implementation of HotSwap, unlike some other approaches, is entirely transparent, without modification to the operating system kernel or the source code of the serverless function. In our tests, loading libraries such as “sklearn + pandas”, “NumPy + keras”, and “NumPy + torch” through HotSwap was up to 2.2x, 3.2x, and 2.5x faster, respectively, compared to traditional AWS Lambda startup processes. Furthermore, experiments in Section IV-E using traces from Azure [11] show that HotSwap can save 88.4% of memory space compared to a function-specific approach, PreBaking [13, 14]. This efficiency is achieved when warming up 10 different serverless functions that share the same dependency.
@@
    This paper proposes HotSwap, a novel cold-start optimization targeting skewed, long-tail serverless workloads, which are common in commercial clouds like Azure.
@@
    HotSwap introduces a new solution to the dependency-loading issue, where the provider stores middleware once in a separate dependency image, independent of the entire container. A single pre-warmed dependency image can be shared across different function instances.
@@
    HotSwap was evaluated for its ability to reduce provider-side costs (CPU usage, cache, RAM) during cold starts using a real-world workload from Azure [11], rather than a synthetic workload.
The rest of this paper is organized as: Section II presents the background and motivation of our paper; Section III presents the design and implementation details of HotSwap; In Section  IV, we evaluate the performance of HotSwap. We then discuss related work in Section V, and conclude in Section VII.
Figure 2: The process of serverless cold starts in the current production cloud
II Background and Motivation
This section introduces: the background for serverless computing and cold starts (Section II-A); the need for HotSwap (Section II-B); and a performance bottleneck in a commercial cloud platform such as AWS Lambda (Section II-C).
II-A Cold Start in Serverless Computing
@@
In Figure 2, while the specifics may differ across platforms, most serverless architectures consist of two core components: a Worker Manager and a Resource Orchestration unit. The Worker Manager responds to a request for capacity from the serverless front end. The Worker Manager acts as a stateful load balancer by continuously monitoring the capacity of each unique function on the platform. If sufficient capacity exists (e.g., idle function instances are available), the Worker Manager directs the front end to route the request to an idle instance. Conversely, if resources are unavailable (triggering a cold start), it instructs the Resource Orchestration unit to provision a new function instance with adequate resources.
Recall also that the user has uploaded the container function to the provider’s Container Registry. Next, the provider will download the user container from the Container Registry and start it. The container initiates the serverless runtime to boot the entire serverless function. Then, the serverless runtime starts a specific function runtime (e.g., Python), and the program loads the necessary software dependencies (e.g., NumPy) and the user-specific handler code. The handler code is then ready to handle a new function instance.
To minimize the cold-start overhead, one needs to skip or expedite one or more steps. When designing a solution for a specific serverless function, all function-specific steps can be considered. Previous work has designed end-to-end, function-specific optimization for targeted serverless functions and achieved substantial reductions in cold-start latency. However, when performing cross-function optimization for the entire system, one can only optimize steps that are identical or similar across all serverless functions, such as container storage and loading.
In this paper, we explore how runtime booting and dependency loading can benefit from cross-function optimization, as they are common processes shared by multiple serverless functions. Thus, we propose HotSwap as a novel cross-function optimization aimed at reducing the time required for runtime booting and dependency loading.
II-B Impact of Invocation Patterns on Cold Starts
In this section, we aim to address two questions: 1) How frequently does a cold start occur for a serverless function? and 2) How does a skewed workload impact the optimization choices of the cloud provider?
@@ II-B Impact of Invocation Patterns on Cold Starts
To address question 1, we model the occurrence of cold starts using an exponential distribution. The frequency of cold starts for a serverless function depends on its invocation pattern. For extreme cases, a function that is never invoked or invoked very frequently (e.g., every second) would not experience cold starts. Let $\lambda$ denote the invocation rate (calls per minute), and $T$ denote the keep-alive time. The probability of no invocation occurring during the keep-alive time follows:
@@
$P_{no\_inv}$ represents the probability that no invocation happens during the keep-alive timeout, leading to a potential cold start.
@@
Figure 1 shows the relation of $E_{cs}$ versus $\lambda$ when $T=15\;min$ and $D=1440\;min$ (a day). Only functions with invocation rates around $1/T$ have a large number of expected cold starts, functions that are less or more frequently invoked have fewer cold starts. Cloud providers like Azure [2] and AWS [1] configure dynamic keep-alive time from $5$ to $30$ minutes. The different keep-alive time changes the position of the maximum point, but the shapes are similar.
To answer question 2, we draw the density histogram of the actual function distribution in the same Figure 1. For function-specific tuning, assuming the cost of optimizing a function is $c$ and the benefits of an optimized cold start are $w$. Functions that are worth function-specific tuning must satisfy the following:
@@
However, from the histogram, the serverless workload in production can be extremely skewed. More than $50\%$ of the functions have a very low invocation rate (less than 0.001 calls per minute) and $E_{cs}$, which experience less than $1.4$ times of cold starts per day. Thus, those functions do not satisfy Equation 3, which means function-specific tuning is not economical for them.
For functions whose invocation rates do not meet the tuning criterion in Equation 3, only cross-function optimizations are feasible. One successful example is AWS Lambda [22], which employs a three-tier caching system and container deduplication to load a 10 GB serverless container in just 50 ms. Such optimizations are applicable to all serverless functions, regardless of invocation frequency.
In this paper, we aim for HotSwap to further enhance cross-function optimization, benefiting serverless functions where function-specific tuning is not economically viable.
II-C Performance Analysis of the Cold Start for Current AWS Lambda Functions